There are 1680 students who attend Central High School. At a recent football game at the school, 40% of the male students and 25% of the female students were in attendance. The total number of students attending the ball game from Central High was 560. How many males and how many females attend this high school?
I first subtracted 100-40 to get 60% and 100-25 to get 75% to multiply by 1680 and used .40*560 and .25*560= I cannot solve the problem. I do not know what I am doing wrong. Please help!
Thanks

Answer by ptaylor(2198)

At least you are givinig it a shot. Here's my approach. I would have thought the numbers of boys and girls would come out in whole numbers but they didn't.
Let x=number of male students
Then 1680-x=number of female students
Now we are told that 40% of the male students (0.40x) plus 25% of the female students (0.25(1680-x)) were in attendence and that amounted to 560 students. So our equation to solve is:
0.40x+0.25(1680-x)=560 get rid of parens
0.40x+420-0.25x=560 subtract 420 from both sides
0.40x+420-420-0.25x=560-420 collect like terms
0.15x=140 divide both sides by 0.15
x=933.3333 or 933----------number of boys
1680-933=747--------------------------------number of girls
0.40(933)+0.25(747)=560
373+187=560
560=560

Hope this helps---ptaylor